# Class Unbiasing for Generalization in Medical Diagnosis

### v1

@@ 4.6.2 Imaging Datasets
BreastMNIST. BreastMNIST contains 780 breast ultrasound images with a resolution of 28 × 28 pixels. It is designed for binary classification task where breast ultrasound images are categorized into “positive” (normal and benign) and “negative” (malignant).
RetinaMNIST. RetinaMNIST is designed for ordinal regression to predict the severity level of diabetic retinopathy, a diabetes-related eye disease, using retina fundus images. The labels correspond to five ordered levels of severity (0-4). We modified the dataset for binary classification by assigning levels 0 and 1 to the negative class, and levels 2-4 to the positive class.
@@ 4.9 Optimization
5 Data Availability
The datasets used in this paper are publicly available or available upon request. DAIC-WOZ is available at https://dcapswoz.ict.usc.edu/. MODMA is available at https://modma.lzu.edu.cn/data/index/. ADReSS is available at https://talkbank.org/dementia/. MedMNIST is availabl at https://zenodo.org/records/10519652.
6 Code Availability
Codes for this work are available at https://github.com/zuo-ls/cls-unbias. All experiments and implementation details are described in the Methods section.

### v2

@@ 5.1.2 Imaging Datasets
BreastMNIST
BreastMNIST contains 780 breast ultrasound images with a resolution of 28 × 28 pixels. It is designed for binary classification task where breast ultrasound images are categorized into “positive” (normal and benign) and “negative” (malignant).
RetinaMNIST
RetinaMNIST is designed for ordinal regression to predict the severity level of diabetic retinopathy, a diabetes-related eye disease, using retina fundus images. The labels correspond to five ordered levels of severity (0-4). We modified the dataset for binary classification by assigning levels 0 and 1 to the negative class, and levels 2-4 to the positive class.
@@ 5.4 Optimization
6 Results
6.1 Main Results
The main results across five datasets are presented in Table 3. Overall, the proposed class unbiased model consistently outperforms the baselines (ERM & ERM (cls-w)), across all datasets.
6.2 Ablation Study
This section investigates the effect of different losses on the DAIC-WOZ dataset, with the results shown in Table 4. In the experiments, the model trained by ERM fails to learn meaningful patterns due to the severe class imbalance in the DAIC-WOZ dataset. Specifically, the model collapses into always predicting the majority (negative) class, thereby effectively minimizing the cross-entropy loss and achieving around 70% accuracy on the training set. Apparently, this seemingly decent performance is an illusion, as the model fails to detect the minority (positive) class. While the variants of ERM partially alleviate this issue—either through a class-weighting strategy (ERM (cls-w)) or by placing equal emphasis on all classes (ERM (per-cls))—the overall performance remains poor, indicating that the model is still negatively affected by class imbalance and potential class-feature bias.
Solely using $\mathcal{L}_{\text{g-dro}}$ (Row 3) is a better approach than using the variants of ERM (ERM (per-cls) & ERM (cls-w)) to handle the class-imbalance problem in the DAIC-WOZ dataset, as the former promotes the exploration of under-represented classes to reduce class loss inequality. However, it fails to handle class-feature bias, because its strategy is to maximize exposure to the minority class rather than minimize reliance on biased class-specific features. In contrast, the class-unbiased models (Rows 4 & 5) substantially improve performance by reducing class loss inequality, thereby mitigating both class imbalance and class-feature bias through explicitly promoting class-wise loss equality, which leads to more generalizable and reliable performance.
Comparing Row 4 and Row 5 of Table 4 further reveals additional improvement brought by incorporating $\mathcal{L}_{\text{g-dro}}$, demonstrating its effectiveness in helping the class-wise inequality loss achieve its intended effect. This is particularly evident on the DAIC-WOZ dataset, where severe class imbalance skews the model’s estimation of $P(Y|X)$, thereby limiting the effectiveness of $\mathcal{L}_{\text{cls-ineq}}$. By introducing $\mathcal{L}_{\text{g-dro}}$, we mitigate the distortion in the estimated $P(Y|X)$ caused by class imbalance, improving the overall performance.
DAIC-WOZ MODMA ADReSS BreastMNIST RetinaMNIST
[…]
Table 4: Ablation study on DAIC-WOZ.
Row Method Valid Test
pos-F1 neg-F1 MF1 Acc pos-F1 neg-F1 MF1 Acc
[…]
6.3 Case Study on Balanced and Imbalanced Datasets.
(a) $\mathcal{L_{\text{cls-ineq}}}$
(b) $\mathcal{L}^{\text{pos}}$ and $\mathcal{L}^{\text{neg}}$.
(c) Classification loss $\mathcal{L}_{\text{ERM}}$ on validation set.
(d) MF1 on validation set.
(a) $\mathcal{L_{\text{cls-ineq}}}$
(b) $\mathcal{L}^{\text{pos}}$ and $\mathcal{L}^{\text{neg}}$.
(c) Classification loss $\mathcal{L}_{\text{ERM}}$ on validation set.
(d) MF1 on validation set.
This section empirically investigates the effect of the proposed class-unbiased approach on balanced and imbalanced datasets. Specifically, we resampled the RetinaMNIST dataset to construct two settings: a balanced dataset with 100 samples (50% positive) and an imbalanced dataset with 100 samples (25% positive). The experimental results are shown in Table 5. The training processes on both balanced and imbalanced datasets are shown in Figures 1 and 2, respectively.
Row Methods cls-balanced MF1 Acc
[…]
7 Discussions & Conclusions
From binary classification to multi-class classification. Extending Cls-unbias from binary classification to multi-class classification is feasible. Similar to [24], one possible strategy is to reduce the variance of the class-wise losses. Encouraging class-wise loss equality remains effective when the number of classes is small. Determining the number of classes for which the multi-class extension of Cls-unbias remains valid can be left to future work.
Finally, we emphasize that Cls-unbias remains highly useful and important in medical diagnosis settings, i.e., binary or few-class classification, where strict feature selection and extraction are critical, as classification failures can lead to severe consequences, including misdiagnosis, delayed treatment, or inadequate monitoring. Moreover, misdiagnosis caused by class-feature bias may result in potential ethical harms. For instance, as discussed in the Section 1, using BMI as a dominant predictive feature for diabetes may lead the model to incorrectly associate high BMI with the disease, reinforcing harmful stereotypes and potentially marginalizing individuals with atypical clinical presentations. Such biased representations can exacerbate healthcare disparities and erode trust in AI-assisted medical decisions.